OpenStudy (anonymous):

Trig: Section - Double Angle Formulas Simplify: 1.sin(pi/12)*cos(pi/12) 2. tan(22.5)/1-tan^2(22.5) | "22.5 is in degrees" Help me please

OpenStudy (alexwee123):

http://www.freemathhelp.com/trig-double-angles.html all the formulas are in there

OpenStudy (anonymous):

I know the formulas but in problem number 1, there is no "2" multiplying sine or cosine?

OpenStudy (anonymous):

then it is half of what you want

OpenStudy (anonymous):

\[\sin(2\theta)=2\sin(\theta)\cos(\theta)\] \[\frac{1}{2}\sin(2\theta)=\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

so \[\frac{1}{2}\sin(\frac{\pi}{6})=\sin(\frac{\pi}{12})\cos(\frac{\pi}{12})\]